40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 8 11:22:35 2014, steve, Summary, SUS, oplev laser summary 
    Reply  Mon Aug 11 10:23:39 2014, steve, Summary, SUS, oplev laser summary updated 
       Reply  Thu Sep 18 10:08:07 2014, steve, Summary, SUS, oplev laser summary updated 
          Reply  Thu Mar 19 17:11:32 2015, steve, Summary, SUS, oplev laser summary updated 
          Reply  Fri Nov 20 09:45:39 2015, steve, Summary, SUS, oplev laser summary  oplev_summery.png
             Reply  Mon May 23 10:55:49 2016, steve, Summary, SUS, ITMX oplev laser replaced ITMXoplev.png
          Reply  Mon May 23 15:51:32 2016, steve, Summary, SUS, oplev laser summary updated oplSum.png
             Reply  Tue May 24 10:21:36 2016, ericq, Summary, SUS, ITMX Oplev loops 
             Reply  Wed Jul 27 11:40:48 2016, steve, Summary, SUS, oplev laser summary updated 
                Reply  Wed Jan 25 10:46:58 2017, steve, Summary, SUS, oplev laser summary updated oplev_sums.png
                   Reply  Wed Feb 8 10:16:39 2017, steve, Update, SUS, clipping ITMX oplev ITMX_oplev_clipping.jpgITMX_clipping.jpg
                      Reply  Fri May 18 07:41:01 2018, steve, Update, SUS, clipping ITMX oplev 
                   Reply  Mon Apr 10 09:58:35 2017, Steve, Summary, SUS, oplev laser summary updated 
                      Reply  Wed Jul 19 12:47:43 2017, Steve, Summary, SUS, oplev laser summary updated 
                         Reply  Wed Jul 19 18:24:15 2017, rana, Summary, SUS, oplev laser summary updated 
                            Reply  Fri Jul 21 08:05:07 2017, Steve, Summary, SUS, red, blue, green laser diodes ordered 
                Reply  Fri Mar 17 15:30:39 2017, Steve, Summary, SUS, oplev laser summary updated oplev_sums.png
Message ID: 12933     Entry time: Mon Apr 10 09:58:35 2017     In reply to: 12753     Reply to this: 13126
Author: Steve 
Type: Summary 
Category: SUS 
Subject: oplev laser summary updated 

 

Quote:

                    Oct.  5, 2015              ETMY He/Ne replaced by 1103P, sr P919645,  made Dec 2014, after 2 years

                   Jan. 24, 2017              ETMY He/Ne replaced by 1103P,  sr P947049,  made Apr 2016,  after 477 hrs running hot

    Jan. 26,  2017              RIN test stared with P947034, made Apr. 2016  

    Apr.  10,  2017              purchased two 1103P from Edmund:  sr P964438 & sr P964431, made 02/2017

   

ELOG V3.1.3-